OpenStudy (anonymous):

2. What is the solution to the system of equations? 3y = 4x + 7 -4x – 4y = 7 A. (-4, -3) B. (-3, -4) C. (4, 3) D. (3, 4)

OpenStudy (anonymous):

One look at your system, and I can see that y has to be zero... are you sure you didn't make a typo? ^.^

OpenStudy (anonymous):

@PeterPan thats what the equation is 3y = 4x + 7 -4x – 4y = 7

OpenStudy (anonymous):

Well, if you rearrange, you'll find -4y = 4x + 7 3y = 4x + 7 If you subtract the two equations -7y = 0 y = 0

OpenStudy (anonymous):

Gotcha what would step two be and I see the 7s cancel out

OpenStudy (anonymous):

Not just the 7's the 4x's too

OpenStudy (anonymous):

What would X be though? 3?

OpenStudy (anonymous):

actually -7/4

OpenStudy (anonymous):

4 *

OpenStudy (anonymous):

Wouldnt that have to make 3 or 4 for x

OpenStudy (anonymous):

If you're not convinced, try substituting the choices, see if any of them fit the system :)

OpenStudy (anonymous):

so would it be c @PeterPan

OpenStudy (anonymous):

Did you check?

OpenStudy (anonymous):

i tried plugging them in but it got confusing and c was the closest answer

OpenStudy (anonymous):

the 7/4 for x is messing me

OpenStudy (anonymous):

:) In Math... there are millions of close answers, and everyone of them is just as wrong as the other :)

OpenStudy (anonymous):

so 4x and -4x would be -4 and for x 3

OpenStudy (anonymous):

4 x + 3 y = 7 .............1 -4 x -4 y = 28 .............2 Eliminate y multiply (1)by 4 Multiply (2) by 3 -16 x 12 y = 28 -12 x -12 y = 84 Add the two equations -28 x = 112 / -28 x = -4 plug value of x in (1) -4 x + 3 y = 7 16 + 3 y = 7 3 y = 7 -16 3 y = -9 y = -3

OpenStudy (anonymous):

@autumnstamm THANK YOU For showing that!!!!

OpenStudy (anonymous):

you saved me i gotta do 2 more like this the steps are gonna help

OpenStudy (anonymous):

You're welcome I thought the steps would help more, more understanding.